OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST ANSWER Which of these nonterminating decimals can be converted into a rational number? Answer choices: 1. 0.874387438743... 2. 0.0000100020003... 3. 5.891420975... 4. 10.074.1259

OpenStudy (anonymous):

Is the answer #1?

OpenStudy (amoodarya):

1

OpenStudy (anonymous):

Thank you!!

OpenStudy (amoodarya):

0.874387438743...=A THEN MULTIPLY BY 10000 10000A=8743.874387438743... SO 10000A-A=8743.874387438743...-0.874387438743... 9999A=8743

OpenStudy (amoodarya):

A=8743/9999 THEN IF POSSIBLE SIMPLIFY